1. A point charge q is placed at a distance r from center of a conducting neutral sphere of radius R (>R). The potential at any point P inside the sphere at a distance r, from point charge due to induced charge of the sphere is given by


1
Expert's answer
2021-06-17T15:07:22-0400

he potential at any point P inside the sphere at a distance r, from point charge due to induced charge of the sphere is given by

"V=\frac{kq}{|\bold{r-r'}|}-\frac{kq}{|\bold{r-\frac{R^2}{r'^2}r'}|}\frac{R}{r'}"